OpenStudy (anonymous):

Use substitution to evaluate the integral:

OpenStudy (anonymous):

\[\int\limits_{0}^{\pi/2} sinxe^(cosx)dx \]

OpenStudy (anonymous):

it is e^(cosx) I'm bad with the equation tool lol

OpenStudy (anonymous):

\[\int_0^{\pi/2}\sin xe^{\cos x}\,dx\] Setting \(u=\cos x\), you have \(-du=\sin x\,dx\), and so \[\int_0^{\pi/2}\sin xe^{\cos x}\,dx=-\int_1^0e^u\,du=\int_0^1e^u\,du\]

OpenStudy (anonymous):

Then what do we do? Substitute back in?

OpenStudy (anonymous):

Or take integral. Confused

OpenStudy (anonymous):

At this point you can simply integrate. \(\int e^u\,du=e^u+C\) It seems to me that the change of variables might have confused you? The first integral is with respect to \(x\) over the interval \(\left[0,\dfrac{\pi}{2}\right]\). When you change the variable, you also have to change the limits. With \(u=\cos x\), for \(x=0\) we get \(u=\cos0=1\), and for \(x=\dfrac{\pi}{2}\) we get \(u=\cos\dfrac{\pi}{2}=0\). The negative sign accounts for the differential, \(du=-\sin x\,dx\). Then I use the property that \(\displaystyle -\int_b^a=\int_a^b\) Now with \[\int_0^1e^u\,du\] you can simply find the antiderivative and apply the fundamental theorem of calculus.

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

You're welcome!